OpenStudy (anonymous):

WILL FAN & GIVE MEDAL Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form. x – 4 = –1(y – 2) x – 2 = –1(y – 4) y – 2 = –1(x – 4) y – 4 = –1(x – 2)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)